OpenStudy (anonymous):

A trapezoid has area 50 in^2. The the bases are 5 in and 15 in. What is the height of the trapezoid?

OpenStudy (amistre64):

still gonna be 5

OpenStudy (anonymous):

but what is the formula?

OpenStudy (amistre64):

(base+base)h ------------ = area 2

OpenStudy (anonymous):

isnt is 1/2(b1+b2)(h)?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

which is what I wrote

OpenStudy (anonymous):

The formula for the area of trapezoid is 1/2(b1+b2)*h A=50 So, we are going to substitute these values to its formula. 50=1/2(5+15)h 100=20h h=5 inches

OpenStudy (anonymous):

but then what do you substitute in for h?

OpenStudy (amistre64):

h is what you solve for...

OpenStudy (amistre64):

h = area(2) ------- (b+b)

OpenStudy (anonymous):

thankss guyss! <3

OpenStudy (anonymous):

you are welcome.